solve for y -8x-8y=-64
8x-8y=-96

Answer by Alan3354(21554)

-8x-8y=-64 --> x+y = 8
8x-8y=-96 --> x-y = -12
-----------
x+y = 8
x-y = -12
----------- Add
2x = -4
x = -2
---------
y = 10